Ken invests $2000 into two separate accounts. One earns 8% interest, the other earns 4.5% interest. If after one year, the total interest from the two accounts is $137.25, how much was invested in each account?
Answer by macston(5194)

.
E=amount at 8%; F=amount at 4.5%
.
E+F=$2000
F=$2000-E Use to substitute for F.
.
0.045F+0.08E=$137.25 Substitute for F from above.
0.045($2000-E)+0.08E=$137.25
$90+0.045E+0.08E=$137.25
0.035E=$47.25
E=$1350 ANSWER 1: $1350 was deposited at 8%.
F=$2000-E=$2000-$1350=$650 ANSWER 2: $650 was deposited at 4.5%.
.
CHECK:
.
0.045F+0.08E=$137.25
0.045($650)+0.08($1350)=$137.25
$29.25+$108=$137.25
$137.25=$137.25